# MC-NeRF: Multi-Camera Neural Radiance Fields for Multi-Camera Image Acquisition Systems

### v2

@@ Abstract
Neural Radiance Fields (NeRF) employ multi-view images for 3D scene representation and have shown remarkable performance. As one of the primary sources of multi-view images, multi-camera systems encounter challenges such as varying intrinsic parameters and frequent pose changes. Most previous NeRF-based methods often assume a global unique camera and seldom consider scenarios with multiple cameras. Besides, some pose-robust methods still remain susceptible to suboptimal solutions when poses are poor initialized. In this paper, we propose MC-NeRF, a method can jointly optimize both intrinsic and extrinsic parameters for bundle-adjusting Neural Radiance Fields. Firstly, we conduct a theoretical analysis to tackle the degenerate case and coupling issue that arise from the joint optimization between intrinsic and extrinsic parameters. Secondly, based on the proposed solutions, we introduce an efficient calibration image acquisition scheme for multi-camera systems, including the design of calibration object. Lastly, we present a global end-to-end network with training sequence that enables the regression of intrinsic and extrinsic parameters, along with the rendering network. Moreover, most existing datasets are designed for unique camera, we create a new dataset that includes four different styles of multi-camera acquisition systems, allowing readers to generate custom datasets. Experiments confirm the effectiveness of our method when each image corresponds to different camera parameters. Specifically, we adopt up to 110 images with 110 different intrinsic and extrinsic parameters, to achieve 3D scene representation without providing initial poses. The Code and supplementary materials are available at https://in2-viaun.github.io/MC-NeRF/
@@
NeRF, Multi-camera system, intrinsic parameters regression, poses regression, 3D scene representation, volume rendering.
@@ 1   Introduction
Visual 3D reconstruction assists autonomous driving, multi-robot navigation, and security surveillance in understanding their environment. These scenes often contain a large number of cameras, which can provide abundant visual information, especially in the form of multi-view perspectives. Recently, Neural Radiance Field (NeRF)[1] demonstrated high quality 3D scene representation capabilities. The essence of NeRF is to learn and reconstruct spatial information by utilizing cross light constraints generated from multi-view images. This information is stored within a neural network to achieve the implicit representation of the scene.
Figure 1: We propose MC-NeRF, which can jointly optimize both intrinsic and extrinsic parameters for bundle-adjusting Neural Radiance Fields. The challenge introduced by multi-camera images is that each image corresponds to different intrinsic and extrinsic parameters, which breaks the original NeRF assumption of a globally unique camera.
@@
An important precondition for NeRF series methods is the acquisition of images from different views. These images often need to be captured using global unique cameras, ensuring uniform camera intrinsic parameters across all images. Specifically, widely utilized NeRF datasets, like Synthesis[1], LLFF[2], NSVF[3], and Mip-NeRF 360[4], employ unique camera intrinsic parameters for each scene, ensuring the ray distribution model is fixed in rendering process. Furthermore, some classical 3D reconstruction approaches also adopted the global unique camera assumption. For instance, Structure from Motion (SFM)[5] assumes the same intrinsic parameters for each image to estimate the poses of individual images. The DTU dataset[6] utilized by MVSNet[7] employs a unique camera for each distinct scene. Based on this unique camera condition, MVSNet constructs a differentiable homography matrix to create a cost volume.
However, in practical scenarios, ensuring that the multi-view images are captured by global unique cameras is not always feasible. A typical application is the multi-camera image acquisition system. We believe the following concerns related to multi-camera acquisition systems: Firstly, each image in the multi-view data captured by the system corresponds to different intrinsic and extrinsic parameters. To establish the ray distribution model within the NeRF series methods, accurate parameters for each image are required. But the calibration process for multi-camera systems can be notably time-consuming. Secondly, the multiple intrinsic matrices invalidates the assumption of a global unique camera, which affects the reliability of methods such as COLMAP[5] used for estimating extrinsic parameters. Besides, frequent adjustments to cameras, like poses and quantities, even camera lens, are common operations in multi-camera acquisition systems, which is accompanied with system vibrations and random camera shifts. These further increase the frequency of system recalibration. In summary, exploring 3D reconstruction methods suitable for multi-camera systems is necessary.
Focusing on the above problems, we propose Multi-Cam NeRF (MC-NeRF), a 3D reconstruction method designed for multi-camera acquisition systems. The method has the capability to directly perform reconstruction using mixed images captured by different cameras, and each image can correspond to a completely different camera intrinsic parameters without requiring any camera extrinsic parameters information. In addition, MC-NeRF achieves end-to-end training and global optimization. During the training process, the intrinsic and extrinsic parameters of each camera and the final NeRF model can be obtained simultaneously. The optimized intrinsic and extrinsic parameters can be stored as calibration parameters, and in this case we can directly use the NeRF series methods for reconstruction, skipping the optimization process of camera parameters. When cameras in the system change (including intrinsic and extrinsic parameters), it only needs to perform global optimization again to obtain new system calibration parameters.
It’s worth noting that our method also has some limitations. Because camera intrinsic parameters can reflect the true scale of images, it is difficult to acquire individual intrinsic parameters for each camera without a standard scale reference. This requires importing additional information for scale calibration. Additionally, during the intrinsic parameter regression process, we encountered a degenerate issue, which prevented us from obtaining valid parameters. This implies that the additional information mentioned above needs to meet certain conditions to avoid degeneration. Moreover, we also found that joint regression of camera parameters leads to parameter coupling, hindering the acquisition of both intrinsic and extrinsic parameters separately. Due to these reasons, our method requires auxiliary calibration images during the optimization process. We also explored the base requirements for calibration and presented a efficient scheme for obtaining these auxiliary images.
@@
As previously discussed, the majority of existing NeRF datasets are generated based on global unique camera, which can not satisfy the requirement of randomized camera parameters with mixed cameras. This has motivated us to propose our own dataset. The MC dataset provides the reader with the flexibility to tailor camera parameters and the number of cameras, allowing for free combinations.
@@
$\bullet$ We address the degenerated cases in intrinsic parameter regression and the coupling problem within joint optimization of intrinsic and extrinsic parameters, and also presented corresponding solutions for these challenges.
@@
$\bullet$ We design a calibration cube for multi-camera image acquisition systems and give an efficient calibration image acquisition scheme.
$\bullet$ We propose a new dataset for multi-camera acquisition systems and provide the source code for dataset generation, enabling readers to create their own datasets freely.
@@ 2   RELATED WORK
Camera Calibration  The camera model demonstrates how points in camera coordinates are projected into pixel coordinates, and camera calibration aims to determine the parameters within this model. Earlier calibration methods observe calibration objects whose geometry in 3D space is known with high precision to calculate the camera parameters, and this efficient method requires only a small number of images to achieve high-performance results[8]. Considering the precision of manufacturing and the cost of calibration objects, Tsai et al. [9] captured images of a 2D planar calibration object from various poses to estimate camera parameters. This method requires providing accurate pose information for each calibration plane. Another widely used approach adopting 2D planar calibration is the chessboard method proposed by Zhang[10]. This technique involves capturing images including a chessboard pattern from different positions to determine the camera parameters. An important precondition of Zhang’s method is that multiple chessboard images need to be captured to ensure the accuracy of the solution. Compared with Tsai’s method, Zhang’s approach eliminates the need to provide the calibrator’s position. In addition to employing known calibration points, there are also some methods relying on detecting specific image features to achieve calibration, like vanishing points[11, 12] and coplanar circles[13].
@@
In recent years, deep learning has provided new inspirations for camera calibration. DeepCalib[14] utilizes a CNN network trained with a vast collection of scene images matched with corresponding camera parameters, and established the relationship between image features and camera parameters. DeepCalib can directly predict calibration results based on the input images. An advantage of this method is that there is no need to provide typical calibration objects. Similar to DeepCalib, Hold-Geoffroy et al.[15] also attempted to predict camera parameters directly from a single image, and they explored the boundaries of human tolerance towards image distortion, and further designed a new perceptual measure for the calibration errors.
@@
Apart from methods designed explicitly to obtain calibration results, some works can indirectly acquire camera parameters. Some learning methods[16, 17, 18, 19, 20, 21, 22, 23] rely on classical multi-view geometry theories, like epipolar constraints, PnP (Perspective-n-Point) algorithms, feature points matching and triangulation, to achieve 3D reconstruction or visual odometry. Although these applications may not design for obtaining camera parameters, they still has the capability to retrieve poses, and the mentioned theories always serve as the foundation for model block design and loss functions design. Additionally, camera pose estimation methods based on NeRF, such as iNeRF[24], have also emerged.
Figure 2: Overall branches of proposed framework. 1) The left side illustrates data acquisition process using a multi-camera system (Room Style). Each camera in the system is fixed in position, and our method requires two packages of calibration images, and one package of the object. 2) The right side displays the loss function branches of the network, consisting of three components. Firstly, it refines the coarse camera intrinsic and extrinsic parameters by utilizing $Pack1$ and $Pack2$. This optimization is executed by the Intrinsic and Extrinsic Reprojection Loss branch. Secondly, with the coarse parameters, the network continues rendering training, incorporating the Photometric Loss branch. This joint optimization can obtain accurate camera parameters and refine the rendering network.
@@
Camera Preconditions of NeRF  Acquiring accurate camera parameters can often be challenging, and in some cases, certain parameters might even be unreachable. When camera parameters are unreliable, adopting classic NeRF-series methods[25, 26, 27, 28, 29, 30, 31, 32, 33] to describe the scene becomes difficult. Some researches[34, 35, 36] employ input images to estimate camera poses. Li et al.[26] used COLMAP to obtain the initial camera poses, which were subsequently used as precondition for NeRF to reconstruct the scene. Instead of using existing methods directly, some other methods[37, 38, 39, 40] can leverage prior knowledge to establish the loss function by introducing geometric constraints with photometric consistency. Specifically, Wang et al.[39] jointly optimized camera parameters and NeRF through photometric reconstruction. Jeong et al[41] design a mixed camera model to achieve the self-calibration, and the process is constrained by geometric and photometric consistency for NeRF training. Lin et al[38] employed a coarse-to-fine auxiliary positional embedding[42, 43, 44] to jointly recover the radiance field and camera pose. NoPe-NeRF[45] uses monocular depth priors to constrain the scene as well as relative pose estimates. Ray-to-ray correspondence losses also can be leveraged to impose constraints on the relative pose estimates by utilizing either keypoint matches or dense correspondences[41, 46]. DBARF[47] adopts low-frequency feature maps to guide the bundle adjustment for generalizable NeRFs[48, 49]. GNeRF[50] introduces a pose estimator to directly estimate camera poses from images. Essentially, these methods combine the training process with camera parameters prediction.
TABLE 1: Description of symbols
Symbol Description
$K_{cam}^{i}$ intrinsic parameters for each camera.
$RT_{local}^{i}$ extrinsic parameters for each camera in $Pack2$
$RT_{glocal}^{i}$ extrinsic parameters for each camera in $Pack1$
( pose for each camera in the multi-camera system)
$Render$ neural radiance field
$W_{local}^{i}$ world coordinates of Apriltag corners in $Pack2$
(ground truth)
$W_{global}^{i}$ world coordinates of Apriltag corners in $Pack1$
(ground truth)
$P_{local}^{i}$ pixel coordinates of Apriltag corners in $Pack2$
(ground truth)
$P_{global}^{i}$ pixel coordinates of Apriltag corners in $Pack1$
(ground truth)
$\tilde{P}_{local}^{i}$ pixel coordinates of Apriltag corners in $Pack2$
$\tilde{P}_{global}^{i}$ pixel coordinates of Apriltag corners in $Pack1$
@@
In summary, NeRF series methods have been extensively explored and significantly improved, yet certain limitations remain. Firstly, previous works often require inputs collected from the same camera or cameras with identical performance, which is difficult to ensure in a multi-camera image acquisition system. Secondly, different intrinsic parameters are rarely considered in prior researches, implying that existing methods aim to standardize all images to a single camera model with uniform parameters.
@@ 3.1   Overview
The main goal of our work is to explore an efficient 3D representation method suitable for multi-camera image acquisition systems. In comparison with previous NeRF series methods, our challenge arises from the unknown intrinsic and extrinsic parameters for each image. The proposed method is illustrated in Fig.2, consisting of three loss functions and can be divided into two branches. In the training stage, we start by jointly training the intrinsic and extrinsic reprojection losses, while masking the rendering branch, to obtain initial intrinsic and extrinsic parameters. Then we employ the initialized parameters to perform optimization of all parameters. The symbol definitions are provided in Table 1.
The reason for employing this strategy is that the inputs for rendering are provided by the intrinsic and extrinsic parameters, as demonstrated in the photometric loss branch. As a downstream component in the sequential flow, the rendering network can be easily influenced by fluctuations in upstream inputs. This implies that training the rendering network may become meaningless when the camera parameters undergo drastic changes. Furthermore, the rendering network adopt coarse to fine structure, which contains a significantly larger number of parameters compared to the intrinsic and extrinsic reprojection loss branches, leading to substantial training time. Therefore, obtaining initialized camera parameters first and subsequently refining them with the rendering network training is an efficient strategy.
Additionally, we employ both the photometric loss branch and the intrinsic reprojection loss branch for joint optimization, with the intrinsic reprojection loss branch serving the purpose of decoupling intrinsic and extrinsic parameters. the issue is detailed discussed in Section 3.4.
3.2   Multi-Camera Intrinsic Parameters Regression
In this section, we first present the theoretical formulas for computing intrinsic parameters. Then, we show how we design learnable parameters. Next, we analyze the degenerate cases during intrinsic parameter optimization, where, in such cases, it becomes challenging to obtain accurate camera intrinsic parameters. Lastly, based on the degenerate cases, we discuss the design scheme and usage of the calibration object.
3.2.1   Theoretical method for intrinsic parameters
The process of computing the camera intrinsic parameters is called calibration. The problem can be described as: given $n$ calibration points $Pi$$(i=0,1,2,...,n)$ in world coordinate system with corresponding pixel coordinates labelled ${{p}_{i}}$ , determining the intrinsic matrix $K$ of the pinhole camera. For the same camera, the projection relationship between $Pi$ and ${{p}_{i}}$ is as follows:
Figure 3: Degenerate case in intrinsic parameters regression. When a single Apriltag is present in the calibration image, obtaining a valid solution is not feasible. However, at least two Apriltags ensures the acquisition of camera intrinsic parameters.
$s{{p}_{i}}=K\left[R\left|T\right.\right]{{P}_{i}}\quad\text{ with }K=\left[\begin{matrix}\alpha&c&{{u}_{0}}\\ 0&\beta&{{v}_{0}}\\ 0&0&1\\ \end{matrix}\right]$ (1)
where $s$ is an arbitrary scale factor, $\left[R\left|T\right.\right]$ are the extrinsic parameters, and $K$ is defined as the camera intrinsic matrix. $({{u}_{0}},{{v}_{0}})$ in $K$ are the coordinates of the principal point, $\alpha$ and $\beta$ are called scale factors in image $u$ and $v$ axes respectively, and $c$ describes the skewness of the two image axes. We define a homography $H$ with size $3\times 4$, which is also called camera matrix, and we have:
$s{{p}_{i}}=H{{P}_{i}}\quad\text{ with }H=K\left[R\left|T\right.\right]$ (2)
with the condition of $n\geq 6$ and ${{\left\|H\right\|}_{F}}=1$, where $F$ is Frobenius norm, $H$ can be solved by SVD decomposition. When $H$ is obtained, $K$ and $R$ can be solved by the first three columns of $H$. Considering that $K$ and $R$ are upper triangular matrix and orthogonal matrix respectively, we can adopt $RQ$ decomposition to get the solutions. Generally, we define the camera coordinate system as same as it defined in OpenCV or Colmap, and the diagonal elements of $K$ are all positive, then we have:
$K=KD\quad\text{ with }D=\left[\begin{matrix}f({{K}_{11}})&0&0\\ 0&f({{K}_{12}})&0\\ 0&0&f({{K}_{13}})\\ \end{matrix}\right]$ (3)
where $f()$ is the sign function $sign()$. Besides, we set the condition of ${{\left\|H\right\|}_{F}}=1$, which is not the real scale of $K$. The final camera intrinsic matrix is defined as $K={K}/{{{K}_{33}}}\;$.
3.2.2   Training Parameters and Degenerated Cases
The core of the network architecture is to reproduce the computing process of Eq.(1). The equation satisfies both the projection mapping and the inverse projection mapping. Projection entails mapping image coordinates ${{p}_{i}}$ to 3D coordinates $Pi$ , while the inverse projection maps 3D coordinates $Pi$ back to image coordinates ${{p}_{i}}$. We choose the inverse projection mapping as the reproduction computation flow because it doesn’t require depth information for precise 2D coordinate retrieval, and the loss function can be implemented in pixel coordinate system.
Assume that the multi-camera system contains $m$ cameras with unknown parameters, and ${{K}_{j}}$$(j=0,1,2,...,m)$ denotes the intrinsic matrix of each camera.
Figure 4: Details of calibration data Acquisition
We decompose ${{K}_{j}}$ into the initialization ${{K}_{j0}}$ and the adjustable weights $\Delta{{K}_{j0}}$, and also assume the axes of the image coordinate system are all perpendicular to each other, which means $c=0$. The matrix:
We define the adjustable weights of extrinsic parameters in the $se(3)$ space and then convert back to the $SE(3)$ space. This approach is taken because representing extrinsic matrix requires only 6 variables in $se(3)$, while it demands 12 variables to regression of $R$ and $T$ in the $SE(3)$ space directly. Let $\left[{{\alpha}_{j0}},{{\alpha}_{j1}},{{\alpha}_{j2}},{{\alpha}_{j3}},{{\alpha}_{j4}},{{\alpha}_{j5}}\right]$$(j=0,1,2,...,m)$ be 6 adjustable weights in the $se(3)$ and the corresponding matrix in the $SE(3)$ is denotes as $\left[{{R}_{j}}\left|{{T}_{j}}\right.\right]$ with size $3\times 4$. For each camera in the system, we have:
$pd\_{{p}_{ij}}={{K}_{j}}\left[{{R}_{j}}\left|{{T}_{j}}\right.\right]{{P}_{ij}}$ (5)
where ${{P}_{ij}}$ is the 3D point used as input and generated by the calibration object, $pd\_{{p}_{ij}}$ is the predicted pixel coordinate. We define the loss function for the intrinsic parameters regression stage as:
$los{{s}_{intr}}=\sum\limits_{i=0,j=0}^{i=n,j=m}{\left\|\frac{gt\_{{p}_{ij}}}{\sqrt{{{h}^{2}}+{{w}^{2}}}}-\frac{pd\_{{p}_{ij}}}{\sqrt{{{h}^{2}}+{{w}^{2}}}}\right\|_{2}^{2}}$ (6)
where $h,w$ represent the height and width of the image, and $gt\_{{p}_{ij}}$ denotes the ground-truth coordinate of ${p}_{ij}$, which can be detected in the images including the calibration object.
It should be noted that the above optimization can not get the correct solutions sometimes, because of the degenerated cases. Specifically, from Eq.(2), we have:
$\left[\begin{matrix}{{u}_{i}}&{{v}_{i}}&1\end{matrix}\right]^{T}=\left[\begin{matrix}h_{1}^{T}\\ h_{2}^{T}\\ h_{3}^{T}\\ \end{matrix}\right]\left[\begin{matrix}{{X}_{i}}&{{Y}_{i}}&{{Z}_{i}}&1\end{matrix}\right]^{T}$ (7)
where ${{h}_{i}}$ has dimensions $4\times 1$, and we convert Eq.(7) to:
$\left\{\begin{matrix}{{u}_{i}}h_{3}^{T}{{P}_{i}}-h_{1}^{T}{{P}_{i}}=0\\ {{v}_{i}}h_{3}^{T}{{P}_{i}}-h_{2}^{T}{{P}_{i}}=0\\ \end{matrix}\right.$ (8)
$\Rightarrow$ $\left\{\begin{matrix}-P_{i}^{T}{{h}_{1}}+0+{{u}_{i}}P_{i}^{T}{{h}_{3}}=0\\ 0-P_{i}^{T}{{h}_{2}}+{{v}_{i}}P_{i}^{T}{{h}_{3}}=0\\ \end{matrix}\right.$
Eq.(8) is a homogeneous system of linear equation consisting of ${{h}_{i}}$. As mentioned earlier, solving this equation requires the condition $n\geq 6$, then we can convert Eq.(8) to:
$Ah=0,\text{ with }$ $A=\left[\begin{matrix}-P_{0}^{T}&0&{{u}_{i}}P_{0}^{T}\\ 0&-P_{0}^{T}&{{v}_{i}}P_{0}^{T}\\ \vdots&\vdots&\vdots\\ -P_{n}^{T}&0&{{u}_{i}}P_{n}^{T}\\ 0&-P_{n}^{T}&{{v}_{i}}P_{n}^{T}\\ \end{matrix}\right]$ (9)

### v4

@@ Abstract
Neural Radiance Fields (NeRF) use multi-view images for 3D scene representation, demonstrating remarkable performance. As one of the primary sources of multi-view images, multi-camera systems encounter challenges such as varying intrinsic parameters and frequent pose changes. Most previous NeRF-based methods assume a unique camera and rarely consider multi-camera scenarios. Besides, some NeRF methods that can optimize intrinsic and extrinsic parameters still remain susceptible to suboptimal solutions when these parameters are poor initialized. In this paper, we propose MC-NeRF, a method for joint optimization of both intrinsic and extrinsic parameters alongside NeRF, allowing individual camera parameters for each image. First, we analyze the coupling issue that arises from the joint optimization between intrinsics and extrinsics, and propose a decoupling constraint utilizing auxiliary images. To further address the degenerate cases in the decoupling process, we introduce an efficient auxiliary image acquisition scheme to mitigate these effects. Furthermore, recognizing that most existing datasets are designed for a unique camera, we provided a new dataset that includes both simulated data and real-world data. Experiments demonstrate the effectiveness of our method in scenarios where each image corresponds to different camera parameters. Specifically, our approach outperforms the baselines favorably in terms of intrinsics estimation, extrinsics estimation, scale estimation, and rendering quality. The Code and supplementary materials are available at https://in2-viaun.github.io/MC-NeRF
@@
NeRF, Multi-camera system, intrinsic parameters estimation, pose estimation, 3D scene representation, volume rendering.
@@ 1   Introduction
Visual 3D reconstruction is pivotal in security surveillance, facial reconstruction, entertainment and robotics. These applications often involve scenes with multi-camera image acquisition systems, providing a wealth of visual information through multi-view perspectives, such as multi-view surveillance systems and facial image capture systems.
Figure 1: We introduce MC-NeRF, which can jointly optimize camera parameters with NeRF. Different from other joint optimization methods, MC-NeRF breaks the assumption of unique camera and does not require providing initial camera parameters.
@@
Recently, Neural Radiance Field (NeRF) [ref1] has shown remarkable capabilities in high-quality 3D scene representation. An important precondition for NeRF series methods is the acquisition of images from different views. These images often need to be captured using a unique camera, ensuring uniform camera intrinsic parameters across all images. Specifically, widely utilized NeRF datasets, like Synthesis [ref1], LLFF [ref2], NSVF [ref3], and Mip-NeRF 360 [ref4], employ unique camera intrinsic parameters for each scene, ensuring the ray distribution model is fixed in rendering process.
However, in multi-camera image acquisition systems, ensuring that all images are captured by a unique camera is not feasible. When using images captured from these systems for reconstruction based on NeRF methods, the following challenges are encountered: Firstly, each image captured by the system corresponds to different intrinsic and extrinsic parameters, which breaks the assumption of camera uniqueness in the NeRF series methods. Secondly, accurately obtaining the parameters of each camera is time-consuming and involves a significant amount of work, and this calibration process must be repeated whenever any camera in the system changes.
A straightforward solution is using Structure-from-Motion (SfM) methods such as COLMAP [ref5] to estimate camera parameters, but these techniques can be brittle and failed, especially with sparse or wide-baseline views [ref55]. Moreover, the camera parameters obtained by COLMAP represent equivalent camera parameters and cannot restore the real-world scale as parameters obtained through calibration.
Another feasible solution involves NeRF methods that jointly optimize camera parameters, such as SC-NeRF [ref41], NeRF--- - [ref39], L2G-NeRF [ref54], and BARF [ref38]. However, these methods still remain susceptible to suboptimal solutions when parameters are poor initialized. Besides, these methods also follow the assumption of a unique camera, which conflicts with the first challenge.
Simply changing the assumption from a unique camera to a multi-camera setup significantly expands the solution space of the joint optimization. This results in the principal points of the intrinsics and the translation items of the extrinsics becoming indistinguishable, preventing them from being accurately estimated. We refer to this as a coupling issue.
Focusing on the challenge, we propose Multi-Camera NeRF (MC-NeRF), a novel 3D reconstruction method for multi-camera image acquisition systems. We establish a global relationship among cameras by introducing additional constraints through the reprojection process. On one hand, the reprojection loss ensures a stable and unique solution space, effectively resolving the coupling issue. On the other hand, given a unique solution for the intrinsics, optimizing the reprojection loss through bundle adjustment (BA) can serve as a method for initializing the extrinsics. Additionally, we also found that not all reprojection processes in a scene can provide effective constraints. We define the cases where they fail as degeneration cases and propose an efficient auxiliary image acquisition scheme to avoid these situations.
@@
As for the datasets, the majority of existing NeRF datasets are generated based on a unique camera, which can not satisfy the requirement of randomized camera parameters with mixed cameras. This has motivated us to propose our own dataset. The MC dataset provides the reader with the flexibility to tailor camera parameters and the number of cameras, allowing for free combinations. Furthermore, to demonstrate the effectiveness of our method in real-world applications, we constructed a multi-camera image acquisition system comprising 88 cameras and provided a dataset from real-world scenarios.
@@
$\bullet$ We address the coupling issue within joint optimization of intrinsic and extrinsic parameters and the degenerated cases in intrinsic parameter estimation.
@@
$\bullet$ We construct a multi-camera image acquisition system comprising 88 cameras and provided a new dataset of real-world scenes captured by this system.
$\bullet$ We provide a new simulated dataset for multi-camera acquisition systems including the source code for dataset generation, enabling readers to create their own datasets freely.
Figure 2: Overview of the proposed method. MC-NeRF utilizes three sets of images as inputs. $Pack1$ is used to unify the world coordinate systems across all cameras and provide initial extrinsics. $Pack2$ imposes intrinsic constraints to address the coupling issue that arises during joint optimization. Multi-view images are scene images, which are used for reconstructing the 3D scene. $Pack1$ and $Pack2$ provide 3D coordinates in space and feature points in the images, which are utilized to establish reprojection loss functions. Then, the intrinsics and extrinsics constrained by the reprojection loss function are used to generate sampling points. These points are subsequently fed into a Multilayer Perceptron (MLP) that employs BARF-based progressive encoding for further training. To ensure both efficiency and convergence, the training sequence must follow the process outlined in Fig.6
@@ 2   RELATED WORK
Camera Calibration  The camera model demonstrates how points in camera coordinates are projected into pixel coordinates, and camera calibration aims to determine the parameters within this model. Earlier calibration methods observe calibration objects whose geometry in 3D space is known with high precision to calculate the camera parameters, and this efficient method requires only a small number of images to achieve high-performance results [ref8]. Considering the precision of manufacturing and the cost of calibration objects, Tsai et al. [ref9] captured images of a 2D planar calibration object from various poses to estimate camera parameters. This method requires providing accurate pose for each calibration plane. Another widely used approach adopting 2D planar calibration is the chessboard method proposed by Zhang [ref10]. This technique involves capturing images including a chessboard pattern from different positions to determine the camera parameters. Compared with Tsai’s method, Zhang’s approach eliminates the need to provide the calibrator’s position. In addition to employing known calibration points, there are also some methods relying on detecting specific features to achieve calibration, like vanishing points [ref11, ref12] and coplanar circles [ref13].
@@
In recent years, deep learning has provided new inspirations for camera calibration. DeepCalib [ref14] utilizes a CNN network trained with a vast collection of scene images matched with corresponding camera parameters, and established the relationship between image features and camera parameters. DeepCalib can directly predict calibration results based on the input images. An advantage of this method is that there is no need to provide typical calibration objects. Similar to DeepCalib, Hold-Geoffroy et al. [ref15] also attempted to predict camera parameters directly from a single image, and they explored the boundaries of human tolerance towards image distortion, and further designed a new perceptual measure for the calibration errors.
@@
Apart from methods designed explicitly to obtain calibration results, some works can indirectly acquire camera parameters. Some learning methods [ref16, ref20] rely on classical multi-view geometry theories, like epipolar constraints, PnP (Perspective-n-Point) algorithms, feature points matching and triangulation, to achieve 3D reconstruction or visual odometry. Although these applications may not design for obtaining camera parameters, they still has the capability to retrieve poses, and the mentioned theories always serve as the foundation for model block design and loss functions design. Additionally, camera pose estimation methods based on NeRF, such as iNeRF [ref24], have also emerged.
@@
Camera Preconditions of NeRF and 3D Gaussian  Acquiring accurate camera parameters can often be challenging, and in some cases, certain parameters might even be unreachable. When camera parameters are unreliable, adopting classic 3D Gaussian [ref56] and NeRF-series methods [ref26, ref31, ref32] to describe the scene becomes difficult. Some researches [ref34, ref35, ref36] employ input images to estimate camera poses. Li et al. [ref26] used COLMAP to obtain the initial camera poses, which were subsequently used as precondition for NeRF to reconstruct the scene. Instead of using existing methods directly, some other methods [ref37, ref40] can leverage prior knowledge to establish the loss function by introducing geometric constraints with photometric consistency. Specifically, Wang et al. [ref39] jointly optimized camera parameters and NeRF through photometric reconstruction. Jeong et al [ref41] design a mixed camera model to achieve the self-calibration, and the process is constrained by geometric and photometric consistency for NeRF training. Lin et al [ref38] employed a coarse-to-fine auxiliary positional embedding [ref42, ref43, ref44] to jointly recover the radiance field and camera pose. NoPe-NeRF [ref45] uses monocular depth priors to constrain the scene as well as relative pose estimates. Ray-to-ray correspondence losses also can be leveraged to impose constraints on the relative pose estimates by utilizing either keypoint matches or dense correspondences [ref41, ref46]. DBARF [ref47] adopts low-frequency feature maps to guide the bundle adjustment for generalizable NeRFs [ref48, ref49]. GNeRF [ref50] introduces a pose estimator to directly estimate camera poses from images. CAMP [ref55] proposes a preconditioner to balance the influence of items in the camera parameters during the joint optimization process, thereby enhancing the quality of the reconstruction. Essentially, these methods combine the training process with camera parameters prediction.
Compared to NeRF, the joint optimization of 3D Gaussians and camera parameters is more complex. Due to the anisotropic characteristics of 3D Gaussian primitives, existing methods such as InstantSplat [ref57], CF-GS [ref58], GS-SLAM [ref59], and SplaTAM [ref60] adopt a fixed camera pose and rotate the scene to obtain the inverse extrinsics, which is then transformed back to the extrinsics. Among these methods, InstantSplat is capable of optimizing both intrinsics and extrinsics simultaneously. It utilizes DUSt3R [ref61] to obtain initial point clouds and camera parameters, and the process of converting intrinsics and extrinsics into learnable parameters is consistent with that of NeRF--- -. Essentially, the idea of these 3D Gaussian methods is to find an optimal combination of camera parameters that renders images with quality consistent with the ground truth images.
@@
In summary, NeRF and 3D Gaussian methods have been extensively explored and significantly improved, yet certain limitations remain. Firstly, previous works often require inputs collected from the same camera or cameras with identical performance, which is difficult to ensure in a multi-camera image acquisition system. Secondly, different intrinsic parameters are rarely considered in prior researches, implying that existing methods aim to standardize all images to a single camera model with uniform parameters.
@@ 3.1   Overview
The proposed method, illustrated in Fig.2. consists of three loss functions and is divided into two branches. It utilizes three sets of images as inputs and does not require any camera initialization parameters. The first set $I$ consists of scene or object images captured by a multi-view system, the second set ${P_{g}}$ includes global auxiliary images, and the third set ${L_{g}}$ contains local auxiliary images, as indicated by the green boxes in Fig.2. The method can be described as follows:
$Scene,Intrs,Extrs=MC-NeRF(I,{P_{g}},{L_{g}})$ (1)
$Scene$ represents the reconstructed scene, while $Intrs$ and $Extrs$ denote the intrinsics and extrinsics of the camera corresponding to the current train image. Since the auxiliary images contain real-world size of AprilTags, the reconstructed scene from MC-NeRF is at real-world scale, which differs from the equivalent scene reconstructed using SFM parameters.
In the following sections, we first address the coupling issues of intrinsics and extrinsics in joint optimization and introduce our solution (Sec 3.2). Next, we discuss the failure cases of the solution, referred to as the degenerate case, and propose an efficient auxiliary image acquisition scheme to avoid this case (Sec 3.3). Following this, we present the training details of our method, including parameterization, loss functions, and the training schedule (Sec 3.4). Finally, we introduce the multi-camera system we used and the public dataset we released (Sec 3.5).
3.2   Coupling Issue in Joint Optimization
We unfold this section by explaining the coupling issue (Sec 3.2.1). We then explore the necessary constraints for jointly optimizing camera parameters and NeRF with existing methods (Sec 3.2.2). Following this, we introduce a solution to the coupling issue when each camera has its own individual intrinsics and extrinsics (Sec 3.2.3).
3.2.1   Coupling in camera parameters
During the joint optimization, $MLP$, $[R|T]$, and $K$ are defined as learnable variables. $[R|T]$ is the camera extrinsics, and $K$ is the camera intrinsics. $enc$ represents the position encoding function, and the execution workflow of NeRF can be described as follows:
$RGB,\sigma$ $=MLP\left({enc(\{{S_{n}}\},\{{D_{n}}\})}\right)\quad$ (2)